# Empirical Bernstein and betting confidence intervals for randomized quasi-Monte Carlo

### v1

@@ 1 Introduction
We study the combination of randomized quasi-Monte Carlo (RQMC) integration to estimate an expectation with some non-asymptotic (finite sample valid) methods to get a confidence interval for such an expectation. RQMC has significant accuracy benefits over plain Monte Carlo (that we outline below), but the usual confidence intervals based on RQMC estimates have justifications that are asymptotic in their sample size LEcEtal24a . Empirical Bernstein and some related betting methods (described below) give finite sample assurances for bounded integrands. Specifically, if it is known that the integrand values must lie between 00 and $1$, then those methods can construct a confidence interval that has at least $1-\alpha$ probability to contain the integral’s value. This confidence level holds for any integrand satisfying the given bounds.
The context for our work is as follows. Many scientific problems require the numerical evaluation of multidimensional integrals. When the dimension is high enough then classical tensor product integration rules like those in DavRab84 become too expensive to use. It is then common to use plain Monte Carlo (MC) methods instead. MC methods based on random sampling typically have root mean squared errors (RMSEs) of $O(n^{-1/2})$ given $n$ function evaluations. This rate is slow but it is the same in any dimension. MC also allows one to use statistical methods to quantify the uncertainty in the estimated integral. Quasi-Monte Carlo (QMC) methods (e.g., DicPil10a and Nie92 ) are deterministic sampling strategies that under a bounded variation assumption produce integral estimates with error $\tilde{O}(n^{-1})$ where $\tilde{O}(\cdot)$ means we neglect powers of $\log(n)$. Plain QMC methods are deterministic and so they lose the uncertainty quantification advantage of MC. RQMC methods, surveyed in LEcLem02a , allow one to make independent replicates of a statistically unbiased QMC method to support variance estimates and asymptotic confidence intervals. RQMC methods can also improve on the convergence rate of QMC methods obtaining an RMSE of $\tilde{O}(n^{-3/2})$ under a smoothness assumption Owe97 ; Owe08a on the integrand.
In addition to asymptotic confidence interval methods for RQMC, some other papers have provided finite sample uncertainty quantifications based on additional assumptions. For purely independent and identically distributed (IID) sampling, knowing a bound on the kurtosis facilitates a finite sample confidence interval for the mean of a random variable in terms of the sample mean and the sample standard deviation HicEtal14a . For deterministic QMC sampling of $\boldsymbol{x}\sim\mathbb{U}[0,1]^{d}$ with lattices or digital nets, assuming that the coefficients in a suitable orthogonal decomposition of $f:[0,1]^{d}\to\mathbb{R}$ decay reasonably, one may derive a deterministic error bound on $\mu=\mathbb{E}[f(\boldsymbol{x})]$ HicJim16a ; HicEtal17a ; JimHic16a ; JagSor23a . A different approach is to assume that $f$ is a realization of a Gaussian process and construct a credible interval for $\mathbb{E}[f(\boldsymbol{x})]$. If the covariance kernel and the sampling nodes are well-matched, the computational effort required is nearly $O(n)$ and not the $O(n^{3})$ effort typically required for such credible intervals RatHic19a ; JagHic22a . The GAIL Gail_ug and QMCPy QMCPy2020a libraries implement these finite sample uncertainty quantifications under the assumptions outlined above.
@@
The non-asymptotic confidence intervals that we emphasize are the recently developed hedged betting confidence intervals (HBCI) and some related predictable plug-in empirical Bernstein confidence intervals (EBCI), both from WauRam24a . They are derived from some infinite confidence sequences and we believe that they provide the narrowest confidence intervals among currently available methods.
Section 2 gives our notation and some definitions and background on quasi-Monte Carlo, randomized quasi-Monte Carlo, and the EBCI and HBCI of WauRam24a . In Section 3 we show how splitting $N$ observations into $R=N/n$ independent replicates of $n$ RQMC point sets can give narrower EBCI than MC does for an oracle that knows the RQMC variance at each $n$. The extent of the narrowing depends on the RQMC convergence rate which varies from problem to problem. Perhaps surprisingly, the more effective RQMC is, the smaller the optimal value of $n$ becomes. For an RQMC variance proportional to $n^{-\theta}$, taking $n$ proportional to $N^{1/(\theta+1)}$ gives the narrowest intervals. The resulting interval widths are $\Theta(N^{-\theta/(\theta+1)})$ which can be much wider than the assumed RQMC standard deviation of $\Theta(N^{-\theta/2})$. Section 4 makes some computational investigations. As predicted by the theory the best values of $n$ to use grew slowly with $N$. The empirically best values of $n$ for HBCI were no larger than the oracle values, and sometimes smaller. Section 5 has some final comments and discussion.
@@ 2 Background and notation
We begin with some notation. We use $1\{E\}$ to denote a quantity that equals $1$ when expression $E$ holds and is zero otherwise. The vectors $\boldsymbol{0}$ and $\boldsymbol{1}$ have all components equal to 00 and $1$ respectively with a dimension given by context. We write $a_{n}=O(b_{n})$ as $n\to\infty$ if there exists $C>0$ and $n_{0}<\infty$ with $|a_{n}|\leqslant Cb_{n}$ whenever $n\geqslant n_{0}$. We write $a_{n}=\Omega(b_{n})$ if there exists $C>0$ and $n_{0}<\infty$ with $a_{n}\geqslant Cb_{n}$ whenever $n\geqslant n_{0}$. We write $a_{n}=\Theta(b_{n})$ if $a_{n}=O(b_{n})$ and $b_{n}=O(a_{n})$.
@@ 2.1 Intervals and martingales
A confidence interval is strict when the right hand side above is an equality. An asymptotic confidence interval is a sequence $(A_{n},B_{n})$ of random quantities for which
@@
The confidence sequences from WauRam24a are derived by applying Ville’s inequality vill:1939 to nonnegative supermartingales.
@@
where $t^{1-\alpha}_{(n-1)}$ is the $1-\alpha$ quantile of Student’s $t$ distribution on $n-1$ degrees of freedom. Then $\hat{\mu}\pm st^{1-\alpha/2}/\sqrt{n}$ is an asymptotic $1-\alpha$ confidence interval for $\mu$.
@@ 2.2 QMC and RQMC
Small values of $D_{n}^{*}$ show that all anchored boxes $[\boldsymbol{0},\boldsymbol{a})$ have very nearly the desired proportion of the $n$ points, which is the volume of $[\boldsymbol{0},\boldsymbol{a})$. It is possible to choose points $\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{n}$ so that $D_{n}^{*}=\tilde{O}(n^{-1})$ while plain MC points have $D_{n}^{*}=\tilde{O}(n^{-1/2})$ Nie92 .
The improvement in star discrepancy yields an improvement in integration error via the Koksma-Hlawka inequality Hic04a :
@@
As described above $D_{n}^{*}$ is a measure of how non-uniform the points $\boldsymbol{x}_{i}$ are. The new factor $V_{\mathrm{HK}}(f)$ is the total variation of $f$ in the sense of Hardy and Krause. See Owe05a . It follows from (3) that we can get $|\hat{\mu}-\mu|=\tilde{O}(n^{-1})$ using QMC. We will refer to this as the BVHK rate.
@@
Here $S^{2}/R$ is an unbiased estimate of the MSE of the pooled estimate $\hat{\mu}$. The confidence level $1-\alpha$ is obtained in the limit as $R\to\infty$ naka:tuff:2024 for fixed $n$.
In addition to getting an unbiased estimate of the RQMC variance, randomization can provide some other benefits. First, some RQMC methods can attain better accuracy than plain QMC. Under sufficient smoothness, the scrambled digital sequences in Owe95 attain RMSEs of $\tilde{O}(n^{-3/2})$ Owe97 ; Owe08a . The same holds for the scramblings of Mat98 . We will refer to this as the ‘smooth case’. Higher order digital nets Dic11a can attain even better convergence rates, though the rates are not necessarily evident empirically when $d$ is moderately large, which nuyens2010higher attribute to numerical precision.
Those RQMC methods have further useful properties. First, $\operatorname{var}(\hat{\mu})=o(1/n)$ as $n\to\infty$ for any integrand with $\sigma^{2}<\infty$, and so their efficiency with respect to MC becomes unbounded as $n\to\infty$. Furthermore, the condition of finite variation in the sense of Hardy and Krause can be quite strict. For example, when $d\geqslant 2$, step discontinuities in the integrand $f$ ordinarily make $V_{\mathrm{HK}}(f)$ infinite unless those discontinuities are axis parallel Owe05a . This makes QMC unattractive for estimating integrals of binary valued functions, while RQMC will still attain an RMSE that is $o(n^{-1/2})$.
While there are numerous RQMC methods in use, we will restrict ourselves here to scramblings of digital nets and sequences, such as the ones of Sobol’ Sob67 . Our variance formulas assume that the scramblings are either those of Owe95 or Mat98 . In our computations we used the linear matrix scramble of Mat98 with a digital shift.
While RQMC methods yield very good accuracy, most of the known confidence interval methods for them supply only asymptotic confidence as $R\to\infty$ naka:tuff:2024 . In some very limited settings there are central limit theorems for $\hat{\mu}$ as $n\to\infty$ Loh01 . Having the confidence statements be asymptotic in $R$ is unfortunate since the RMSE of $\hat{\mu}$ is proportional to $R^{-1/2}$ while the RMSE vanishes at a faster rate in $n$.
@@ 2.3 Hoeffding and Empirical Bernstein intervals
There is a theorem of Bahadur and Savage BahSav56 that reveals sharp constraints on our ability to construct nonparametric and non-asymptotic confidence intervals. Here is the description from err4qmc :
They consider a set $\mathcal{F}$ of distributions on $\mathbb{R}$. Letting $\mu(F)$ be $\mathbb{E}(Y)$ when $Y\sim F\in\mathbb{R}$, their conditions are:
@@
    $\mathcal{F}$ is convex: if $F,G\in\mathcal{F}$ and $10<\pi<$ 1, then $\pi F+(1-\pi)G\in\mathcal{F}$.
@@
for $10<\alpha<$ 1 is a $1-\alpha$ confidence interval for $\mu=\mathbb{E}(Y_{i})$. Equation (4) gives the confidence interval of Hoeffding hoef:1963 . The interval is not always nested within $[0,1]$, but we can simply take its intersection with $[0,1]$. That will preserve the confidence level.
While Hoeffding’s interval is a valid confidence interval for the mean of any distribution supported on $[0,1]$, it can be quite conservative for some of them, giving confidence over $1-\alpha$. This raises the possibility of getting coverage $1-\alpha$ from narrower intervals. One approach is to make use of $\sigma^{2}=\operatorname{var}(Y)$ if it is known. Hoeffding’s confidence interval is a consequence of Hoeffding’s exponential probability bound in hoef:1963 . Other exponential probability bounds take advantage of a known value $\sigma^{2}=\operatorname{var}(Y)$ or even a sample estimate of that variance. Those bounds can be used to get improved confidence intervals.
Theorem 3 of Maurer and Pontil maurer2009empirical is that
@@
holds with probability at least $1-\delta$. They call this Bennett’s inequality, because it derives from an inequality in benn:1962 . They do not supply a proof, but it can be proved from Bernstein’s inequality given as Theorem 3 of bouc:lugo:bous:2003 . Taking $\delta=\alpha/2$ we get
@@
In most practical settings where $\mu$ is unknown, $\sigma^{2}$ will also be unknown. Replacing $\sigma^{2}$ by $s^{2}$ in (5) would not always give a valid confidence interval. Theorem 4 of maurer2009empirical shows that for $R\geqslant 2$
@@
is a $1-\alpha$ confidence interval for $\mu$. This is called an empirical Bernstein confidence interval named after an inequality of Bernstein that is similar to Bennett’s. We see that both terms on the right hand side of (6) are larger than the ones in the Bennett interval (5). Their interval has improved asymptotic constants compared to the original empirical Bernstein inequality from audi:muno:szep:2007 . The EBCI and HBCI from WauRam24a provide a further asymptotic improvement.
@@ 2.4 Betting methods
Our interest in this problem was sparked in part by some recent work by WauRam24a on confidence intervals generated by betting arguments. They consider a more general setting of confidence sequences, as in equation (2), that can then be specialized to confidence intervals as needed. For random $Y_{i}\in[0,1]$ they assume that $\mathbb{E}(Y_{i}\!\mid\!Y_{1},\dots,Y_{i-1})=\mu$ for all $i\geqslant 1$ without assuming that the $Y_{i}$ are identically distributed or even that they are independent. For $i=1$, the conditional expectation above is simply $\mathbb{E}(Y_{1})=\mu$.
Suppose that the conditional mean is really $\mu$ and there is a null hypothesis $H_{0}(m)$ that $\mathbb{E}(Y_{i}\!\mid\!Y_{1},\dots,Y_{i-1})=m$. Then if $m\neq\mu$ somebody starting with a stake of \$1.00 can make a series of bets against $H_{0}(m)$ as the $Y_{i}$ are revealed in order and have an expected fortune that grows without bound. Just prior to time $i$, the bettor picks a value $\lambda_{i}\in(-1/(1-m),1/m)$. The bettor’s capital at time $t\geqslant 1$ is
@@
Section 4 of WauRam24a notes that $\mathcal{K}_{t}(\mu)$ is a nonnegative martingale, so that $\mathbb{E}(\mathcal{K}_{t}(\mu))=1$ for all $t\geqslant 1$. There is no way for the bettor to pick bet sizes $\lambda_{i}(\mu)$ to make an expected profit. Then from Ville’s Theorem
@@
If we watch $\mathcal{K}_{t}(\mu)$ indefinitely, then there is at most probability $\alpha$ that it will ever go above $1/\alpha$. As a result we can get a confidence sequence by retaining at time $t$, all the values $m$ for which $\max_{1\leqslant\tau\leqslant t}\mathcal{K}_{\tau}(m)\leqslant 1/\alpha$. This idiom uses a hypothetically infinite ensemble of bettors to do this, but WauRam24a give implementable algorithms for it.
@@
for some $c\in(0,1)$ with $c=1/2$ or $3/4$ given as reasonable defaults. Note that this method only bets that $Y_{i}>m$. Theorem 2 of WauRam24a gives a $1-\alpha$ confidence sequence
@@
These confidence sequences can be specialized to confidence intervals when we have a fixed target sample size $R$ in mind. For that case WauRam24a recommends
@@
The above formula for $\lambda_{i}$ uses the ordering of the data values and puts unequal weight on the $R$ different $Y_{i}$. There is a permutation strategy in WauRam24a to treat the $Y_{i}$ values more symmetrically, but they report that it makes little difference.
@@
where $C_{t}^{\text{PrPl-EB}(R)}$ is the interval in (7) using $\lambda_{i}$ from (8). Equation (17) of WauRam24a gives the scaled half width of these EBCI as
@@
as $R\to\infty$. The corresponding limit for the empirical Bernstein CIs of maurer2009empirical is $\sigma\sqrt{2\log(4/\alpha)}$.
The hedged betting process of WacMenSch08 works as follows. For positive predictable sequences $\lambda_{i}^{+}(m)$ and $\lambda_{i}^{-}(m)$ defined below, they use capital processes
$K_{t}^{+}(m)=\prod_{i=1}^{t}(1+\lambda_{i}^{+}(m)(Y_{i}-m)\quad\text{and}\quad K_{t}^{-}(m)=\prod_{i=1}^{t}(1-\lambda_{i}^{+}(m)(Y_{i}-m)$
@@
for $0\leqslant\theta\leqslant 1$. For any $\theta$, $K_{t}^{\pm}(\mu)$ is a nonnegative martingale hence subject to Ville’s theorem. The supplement to WauRam24a suggests $\theta=1/2$ as a default. If $m\neq\mu$ then $K_{t}^{\pm}(m)$ will tend to grow with $t$. Of course the hedging $\theta$ has nothing to do with the RQMC variance rate $\theta$ and the quantities are distinct enough that no confusion should arise.

### v2

@@ 1 Introduction
We study the combination of randomized quasi-Monte Carlo (RQMC) integration to estimate an expectation with some non-asymptotic (finite sample valid) methods to get a confidence interval for such an expectation. RQMC has significant accuracy benefits over plain Monte Carlo (that we outline below), but the usual confidence intervals based on RQMC estimates have justifications that are asymptotic in their sample size Nakayama and Tuffin (2024). Empirical Bernstein and some related betting methods (described below) give finite sample assurances for bounded integrands. Specifically, if it is known that the integrand values must lie between 0 and $1$, then those methods can construct a confidence interval that has at least $1-\alpha$ probability to contain the integral’s value. This confidence level holds for any integrand satisfying the given bounds.
[…]
In addition to asymptotic confidence interval methods for RQMC, some other papers have provided finite sample uncertainty quantifications based on additional assumptions. For purely independent and identically distributed (IID) sampling, knowing a bound on the kurtosis facilitates a finite sample confidence interval for the mean of a random variable in terms of the sample mean and the sample standard deviation Hickernell et al. (2013). For deterministic QMC sampling of $\boldsymbol{x}\sim\mathbb{U}[0,1]^{d}$ with lattices or digital nets, assuming that the coefficients in a suitable orthogonal decomposition of $f:[0,1]^{d}\to\mathbb{R}$ decay reasonably, one may derive a deterministic error bound on $\mu=\mathbb{E}[f(\boldsymbol{x})]$ Hickernell and Jiménez Rugama (2016); Hickernell et al. (2018); Jiménez Rugama and Hickernell (2016); Sorokin and Jagadeeswaran (2024). A different approach is to assume that $f$ is a realization of a Gaussian process and construct a credible interval for $\mathbb{E}[f(\boldsymbol{x})]$. If the covariance kernel and the sampling nodes are well-matched, the computational effort required is nearly $O(n)$ and not the $O(n^{3})$ effort typically required for such credible intervals Jagadeeswaran and Hickernell (2019, 2022). The GAIL Choi et al. (2015) and QMCPy Choi et al. (2025) libraries implement these finite sample uncertainty quantifications under the assumptions outlined above.
@@
The non-asymptotic confidence intervals that we emphasize are the recently developed hedged betting confidence intervals (HBCI) and some related predictable plug-in empirical Bernstein confidence intervals (EBCI), both from Waubdy-Smith and Ramdas (2024). They are derived from some infinite confidence sequences and we believe that they provide the narrowest confidence intervals among currently available methods.
Section 2 gives our notation and some definitions and background on quasi-Monte Carlo, randomized quasi-Monte Carlo, and the EBCI and HBCI of Waubdy-Smith and Ramdas (2024). In Section 3 we show how splitting $N$ observations into $R=N/n$ independent replicates of $n$ RQMC point sets can give narrower EBCI than MC does for an oracle that knows the RQMC variance at each $n$. The extent of the narrowing depends on the RQMC convergence rate which varies from problem to problem. Perhaps surprisingly, the more effective RQMC is, the smaller the optimal value of $n$ becomes. For an RQMC variance proportional to $n^{-\theta}$, taking $n$ proportional to $N^{1/(\theta+1)}$ gives the narrowest intervals. The resulting interval widths are $\Theta(N^{-\theta/(\theta+1)})$ which can be much wider than the assumed RQMC standard deviation of $\Theta(N^{-\theta/2})$. Section 4 makes some computational investigations. As predicted by the theory the best values of $n$ to use grow slowly with $N$. The empirically best values of $n$ for HBCI are no larger than the oracle values, and sometimes smaller. Section 5 has some final comments and discussion.
@@ 2 Background and notation
We begin with some notation. We use $1\{E\}$ to denote a quantity that equals $1$ when expression $E$ holds and is zero otherwise. The vectors $\boldsymbol{0}$ and $\boldsymbol{1}$ have all components equal to 0 and $1$ respectively with a dimension given by context. We write $a_{n}=O(b_{n})$ as $n\to\infty$ if there exists $C>0$ and $n_{0}<\infty$ with $|a_{n}|\leqslant Cb_{n}$ whenever $n\geqslant n_{0}$. We write $a_{n}=\Omega(b_{n})$ if there exists $C>0$ and $n_{0}<\infty$ with $a_{n}\geqslant Cb_{n}$ whenever $n\geqslant n_{0}$. We write $a_{n}=\Theta(b_{n})$ if $a_{n}=O(b_{n})$ and $b_{n}=O(a_{n})$.
@@ 2.1 Intervals and martingales
A confidence interval is strict when the sides of (1) are equal. An asymptotic confidence interval is a sequence $(A_{n},B_{n})$ of random quantities for which
@@
The confidence sequences from Waubdy-Smith and Ramdas (2024) are derived by applying Ville’s inequality Ville (1939) to nonnegative supermartingales.
@@
where $t^{1-\alpha}_{(n-1)}$ is the $1-\alpha$ quantile of Student’s $t$ distribution on n1n-1 degrees of freedom. Then $\hat{\mu}\pm st^{1-\alpha/2}/\sqrt{n}$ is an asymptotic $1-\alpha$ confidence interval for $\mu$.
@@ 2.2 QMC and RQMC
Small values of $D_{n}^{*}$ show that all anchored boxes $\boldsymbol{0},\boldsymbol{a})$ have very nearly the desired proportion of the $n$ points, which is the volume of $\boldsymbol{0},\boldsymbol{a})$. It is possible to choose points $\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{n}$ so that $D_{n}^{*}=\tilde{O}(n^{-1})$ while plain MC points have $D_{n}^{*}=\tilde{O}(n^{-1/2})$ Niederreiter (1992).
The improvement in star discrepancy yields an improvement in integration error via the Koksma-Hlawka inequality Hickernell (2006):
@@
As described above $D_{n}^{*}$ is a measure of how non-uniform the points $\boldsymbol{x}_{i}$ are. The new factor $V_{\mathrm{HK}}(f)$ is the total variation of $f$ in the sense of Hardy and Krause. See Owen (2005). It follows from (3) that we can get $|\hat{\mu}-\mu|=\tilde{O}(n^{-1})$ using QMC. We will refer to this as the BVHK rate.
@@
Here $S^{2}/R$ is an unbiased estimate of the MSE of the pooled estimate $\hat{\mu}$. The confidence level $1-\alpha$ is obtained in the limit as $R\to\infty$ Nakayama and Tuffin (2024) for fixed $n$. That convergence is generally quicker for more nearly Gaussian $\hat{\mu}_{i}$.
In addition to getting an unbiased estimate of the RQMC variance, randomization can provide some other benefits. First, some RQMC methods can attain better accuracy than plain QMC. Under sufficient smoothness, the scrambled digital sequences in Owen (1995) attain RMSEs of $\tilde{O}(n^{-3/2})$ Owen (1997, 2008). The same holds for the scramblings of Matoušek (1998). We will refer to this as the ‘smooth case’. Higher order digital nets Dick (2011) can attain even better convergence rates, though the rates are not necessarily evident empirically when $d$ is moderately large, which Nuyens and Cools (2010) attribute to numerical precision.
Those RQMC methods have further useful properties. First, $\operatorname{var}(\hat{\mu})=o(1/n)$ as $n\to\infty$ for any integrand with $\sigma^{2}<\infty$, and so their efficiency with respect to MC becomes unbounded as $n\to\infty$. Furthermore, the condition of finite variation in the sense of Hardy and Krause can be quite strict. For example, when $d\geqslant 2$, step discontinuities in the integrand $f$ ordinarily make $V_{\mathrm{HK}}(f)$ infinite unless those discontinuities are axis parallel Owen (2005). This makes QMC unattractive for estimating integrals of binary valued functions, while RQMC will still attain an RMSE that is $o(n^{-1/2})$.
While there are numerous RQMC methods in use, we will restrict ourselves here to scramblings of digital nets and sequences, such as the ones of Sobol’ Sobol’ (1967). Our variance formulas assume that the scramblings are either those of Owen (1995) or Matoušek (1998). In our computations we used the linear matrix scramble of Matoušek (1998) with a digital shift.
While RQMC methods yield very good accuracy, most of the known confidence interval methods for them supply only asymptotic confidence as $R\to\infty$ Nakayama and Tuffin (2024). In some very limited settings there are central limit theorems for $\hat{\mu}$ as $n\to\infty$ Loh (2003). Having the confidence statements be asymptotic in $R$ is unfortunate since the RMSE of $\hat{\mu}$ is proportional to $R^{-1/2}$ while the RMSE vanishes at a faster rate in $n$.
@@ 2.3 Hoeffding and Empirical Bernstein intervals
There is a theorem of Bahadur and Savage Bahadur and Savage (1956) that reveals sharp constraints on our ability to construct nonparametric and non-asymptotic confidence intervals. Here is the description from Owen (2026):
They consider a set $\mathcal{F}$ of distributions on $\mathbb{R}$. Letting $\mu(F)=\mathbb{E}(Y)$ for $Y\sim F\in\mathcal{F}$, their conditions are:
@@
    $\mathcal{F}$ is convex: if $F,G\in\mathcal{F}$ and $0<\pi<1$, then $\pi F+(1-\pi)G\in\mathcal{F}$.
@@
for $0<\alpha<1$ is a $1-\alpha$ confidence interval for $\mu=\mathbb{E}(Y_{i})$. Equation (4) gives the confidence interval of Hoeffding Hoeffding (1963). The interval is not always nested within $[0,1]$, but we can simply take its intersection with $[0,1]$. That will preserve the confidence level.
While Hoeffding’s interval is a valid confidence interval for the mean of any distribution supported on $[0,1]$, it can be quite conservative for some of them, giving confidence over $1-\alpha$. This raises the possibility of getting coverage $1-\alpha$ from narrower intervals. One approach is to make use of $\sigma^{2}=\operatorname{var}(Y)$ if it is known. Hoeffding’s confidence interval is a consequence of Hoeffding’s exponential probability bound in Hoeffding (1963). Other exponential probability bounds take advantage of a known value $\sigma^{2}=\operatorname{var}(Y)$ or even a sample estimate of that variance. Those bounds can be used to get improved confidence intervals.
Theorem 3 of Maurer and Pontil Maurer and Pontil (2009) is that
@@
holds with probability at least $1-\delta$. They call this Bennett’s inequality, because it derives from an inequality in Bennett (1962). They do not supply a proof, but it can be proved from Bernstein’s inequality given as Theorem 3 of Boucheron et al. (2003). Taking $\delta=\alpha/2$ we get
@@
In most practical settings where $\mu$ is unknown, $\sigma^{2}$ will also be unknown. Replacing $\sigma^{2}$ by $s^{2}$ in (5) would not always give a valid confidence interval. Theorem 4 of Maurer and Pontil (2009) shows that for $R\geqslant 2$
@@
is a $1-\alpha$ confidence interval for $\mu$. This is called an empirical Bernstein confidence interval named after an inequality of Bernstein that is similar to Bennett’s. We see that both terms on the right hand side of (6) are larger than the ones in the Bennett interval (5). Their interval has improved asymptotic constants compared to the original empirical Bernstein inequality from Audibert et al. (2007). The EBCI and HBCI from Waubdy-Smith and Ramdas (2024) provide a further asymptotic improvement.
@@ 2.4 Betting methods
Our interest in this problem was sparked in part by some recent work by Waubdy-Smith and Ramdas (2024) on confidence intervals generated by betting arguments. They consider a more general setting of confidence sequences, as in equation (2), that can then be specialized to confidence intervals as needed. For random $Y_{i}\in[0,1]$ they assume that $\mathbb{E}(Y_{i}\!\mid\!Y_{1},\dots,Y_{i-1})=\mu$ for all $i\geqslant 1$ without assuming that the $Y_{i}$ are identically distributed or even that they are independent. For $i=1$, the conditional expectation above is simply $\mathbb{E}(Y_{1})=\mu$.
Suppose that the conditional mean is really $\mu$ and there is a null hypothesis $H_{0}(m)$ that $\mathbb{E}(Y_{i}\!\mid\!Y_{1},\dots,Y_{i-1})=m$. Then if $m\neq\mu$ somebody starting with a stake of \$1.00 can make a series of bets against $H_{0}(m)$ as the $Y_{i}$ are revealed in order and have an expected fortune that grows without bound. Just prior to time $i$, the bettor picks a value $\lambda_{i}=\lambda_{i}(m)\in(-1/(1-m),1/m)$. The bettor’s capital at time $t\geqslant 1$ is
@@
Section 4 of Waubdy-Smith and Ramdas (2024) notes that $\mathcal{K}_{t}(\mu)$ is a nonnegative martingale, so that $\mathbb{E}(\mathcal{K}_{t}(\mu))=1$ for all $t\geqslant 1$. There is no way for the bettor to pick bet sizes $\lambda_{i}(\mu)$ to make an expected profit. Then from Ville’s Theorem
@@
If we watch $\mathcal{K}_{t}(\mu)$ indefinitely, then there is at most probability $\alpha$ that it will ever go above $1/\alpha$. As a result we can get a confidence sequence by retaining at time $t$, all the values $m$ for which $\max_{1\leqslant\tau\leqslant t}\mathcal{K}_{\tau}(m)\leqslant 1/\alpha$. This idiom uses a hypothetically infinite ensemble of bettors to do this, but Waubdy-Smith and Ramdas (2024) give implementable algorithms for it.
@@
for some $c\in(0,1)$ with $c=1/2$ or $3/4$ given as reasonable defaults. Note that this method only bets that $Y_{i}>m$. Theorem 2 of Waubdy-Smith and Ramdas (2024) gives a $1-\alpha$ confidence sequence
@@
These confidence sequences can be specialized to confidence intervals when we have a fixed target sample size $R$ in mind. For that case Waubdy-Smith and Ramdas (2024) recommends
@@
The above formula for $\lambda_{i}$ uses the ordering of the data values and puts unequal weight on the $R$ different $Y_{i}$. There is a permutation strategy in Waubdy-Smith and Ramdas (2024) to treat the $Y_{i}$ values more symmetrically, but they report that it makes little difference.
@@
where $C_{t}^{\text{PrPl-EB}(R)}$ is the interval in (7) using $\lambda_{i}$ from (8). Equation (17) of Waubdy-Smith and Ramdas (2024) gives the scaled half width of these EBCI as
@@
as $R\to\infty$. The corresponding limit for the empirical Bernstein CIs of Maurer and Pontil (2009) is $\sigma\sqrt{2\log(4/\alpha)}$.
The hedged betting process of Wackerly et al. (2008) works as follows. For positive predictable sequences $\lambda_{i}^{+}(m)$ and $\lambda_{i}^{-}(m)$ defined below, they use capital processes
$K_{t}^{+}(m)=\prod_{i=1}^{t}\bigl(1+\lambda_{i}^{+}(m)(Y_{i}-m)\bigr)\quad\text{and}\quad K_{t}^{-}(m)=\prod_{i=1}^{t}\bigl(1-\lambda_{i}^{+}(m)(Y_{i}-m)\bigr)$
@@
for $0\leqslant\theta\leqslant 1$. For any $\theta$, $K_{t}^{\pm}(\mu)$ is a nonnegative martingale hence subject to Ville’s theorem. The supplement to Waubdy-Smith and Ramdas (2024) suggests $\theta=1/2$ as a default. If $m\neq\mu$ then $K_{t}^{\pm}(m)$ will tend to grow with $t$. Of course the hedging $\theta$ has nothing to do with the RQMC variance rate $\theta$ and the quantities are distinct enough that no confusion should arise.